R-388 s/n 193

Received for work 11/05.

Initial survey:

    Generally quite dirty, but mostly just dust buildup, should clean up nicely.  Front panel very good, some scars at the rack-mount holes have been previously touched up.  Has a home-made bottom cover, no top cover, has the original 2 phenolic alignment tools.  The dial drum is not very yellowed, and there are no noticeable scratches or tears noticed.  Has holdowns for the calibrator xtal and the filter cap.  Has an SS replacement plug-in module in place of the 5V4G rectifier tube.  It has original wheels on the back bottom corners of the chassis for easy placement into a cabinet.  There is one possible modification noted underneath, and V-110 & C-205.  No other previous work noted, tho' it's possible the AC line cord (2-wire) has been replaced.  The owner had indicated that there had been some damage to some of the RF/IF coils and/or the fine wiring to them, underneath.  None was immediately evident.  The "canned" electrolytic filter capacitor sections check at 0.29 & 0.32 ohms ESR, good.  Its' date code is 3-57.  

    It is powered up on a variac fairly quickly, as the filter caps ckd OK.  The MC indicator drum is mis-indexed, showing 1 or 2 bands higher than actually received.  The calibrator signal is heard on all bands, although sensitivity is low on bands above 15mc.  The AVC holds sensitivity down more than proper, but it is still low, at least on higher bands.  Audio output may be low.  The PTO is off 6-1/2 kc from end to end.  The xtal calibrator signal can only be trimmed to about 200 cycles off freq. at 15mc.

"before" pix

b4panel1.jpg (236883 bytes) b4panel5.jpg (178081 bytes) b4panel4.jpg (240003 bytes) b4panel3.jpg (246551 bytes) b4panel2.jpg (259466 bytes) b4panel6.jpg (211024 bytes)


b4top1.jpg (233099 bytes) b4bot3.jpg (294315 bytes) b4bot2.jpg (294330 bytes) 

"after" pix

The top(and underneath) is much nicer looking after cleanup.

  R388sn193images/aftTop1.JPG (76614 bytes) R388sn193images/aftTop2.JPG (78554 bytes) R388sn193images/aftTop3.JPG (72235 bytes) 

PTO work

After cleanup, the first thing to be done is the PTO. It's a Collins mod. 70E-15, s/n M-10905. After changing the feedback capacitor and it was setup on the bench, readings were taken every 100kc, and it was found to have 6.52 kc less than 1mc in 10 turns. The compensation coil slug had been previously adjusted to near it's maximum range, and this much drift is probably more than it could compensate for if it hadn't been used before. Just less than one turn was removed from the compensating coil, and more readings were taken. After several iterations of adjusting the compensating slug and taking readings it ended up being 200 cycles short of 1000kc in ten turns, and less than 220 cycles difference between 100kc points on all but one 100kc section, where it was off 420 cycles. The replaced feedback capacitor was found to have approx. 4 megohms of resistance at 300 volts, not terrible, but should have been 100 megohms, or more. The calibration of each band may be shifted by 1kc or so, noticeable when switching between bands, as the individual band crystals may have drifted slightly. This is easily adjusted by the "zero" adjustment, using the crystal calibrator.

R388sn193images/PTO1.JPG (55228 bytes) R388sn193images/PTO6.JPG (70724 bytes) R388sn193images/PTO5.JPG (63135 bytes)

Reassembly and component replacement

All tubes were tested before replacement. 1 6BA6 & 1 6BE6 checked poor and were replaced with new ones. The electrolytic filter capacitor, C-217A&B was "restuffed" with new components. The 2-wire AC line cord was replaced with a new 3-wire one.

The following "bathtub" capacitors were removed and replaced with discrete components mounted on terminal strips:
C-198A&B, 0.1 mf, 3rd IF amp cath. bypass, fil. bypass at T-108
C-205A,B&C, 0.1 mf , back bias rf filter, AVC amp stabilizing, Noise lim. filter
C-215, 20mf, IF output cath. bypass
C-216, 20mf, bias filter
C-223 8mf, B+ isolation
C239 8mf, 150 reg. line filter

A 0.1 capacitor was added on the 150v regulated line at the voltage regulator tube to eliminate hash, and a 1000ohm 2watt resistor added in the voltage dropping line to the VR tube to decrease bias voltage by about 5v. The bias voltage is still lower than original due to higher AC line input voltage. It may be acceptable.

C-204, 100pf, AVC rectifier coupling, a common problem in these receivers, had been replaced by a previous owner. It was checked as OK and retained.

The IF's, RF's and BFO were aligned.

Several relatively minor items still need to be addressed, e.g. the band drum misalignment is partly due to what appears to be improper placement of the drive drum on the switch shaft, and the over-travel stop on the PTO.


The band drum misalignment and insufficient travel has been fixed, as has the PTO overtravel on each end.  The crystal filter phasing has been aligned and is working well.   Some small nicks and scrapes were touched up on the front panel.  Time spent just listening around. 
  One thing I didn't see in the initial evaluation, but noticed later, is that the front panel had been "over-sprayed".  Each area of lettering had been masked with narrow tape before the whole panel was sprayed.  The result is that the wrinkle paint at the lettering is slightly lighter than the rest of the panel.  It doesn't appear that the original wrinkle fiinsh was badly nicked or scraped, and overall it is good looking, from a distance.  Only from up close is the masking effect noticeable, it may not be visible in the pix. 

R388sn193images/done7.JPG (64272 bytes) R388sn193images/done3.JPG (67422 bytes) R388sn193images/done2.JPG (74295 bytes) R388sn193images/done1.JPG (76711 bytes) R388sn193images/done6r.jpg (172411 bytes) R388sn193images/done8.JPG (84388 bytes)

This one's done, after I install the 15 screws in the bottom cover, and the "wheels" on the back corners of the chassis.


Radio to see home page

comments to anchor@ec.rr.com

And remember: "They don't make tubes nowadays like they used to..."











Hit Counter